# Diffusion Model Based Signal Recovery Under 1-Bit Quantization

### v1

@@ Abstract
Diffusion models (DMs) have demonstrated to be powerful priors for signal recovery, but their application to 1-bit quantization tasks, such as 1-bit compressed sensing and logistic regression, remains a challenge. This difficulty stems from the inherent non-linear link function in these tasks, which is either non-differentiable or lacks an explicit characterization. To tackle this issue, we introduce Diff-OneBit, which is a fast and effective DM-based approach for signal recovery under 1-bit quantization. Diff-OneBit addresses the challenge posed by non-differentiable or implicit links functions via leveraging a differentiable surrogate likelihood function to model 1-bit quantization, thereby enabling gradient based iterations. This function is integrated into a flexible plug-and-play framework that decouples the data-fidelity term from the diffusion prior, allowing any pretrained DM to act as a denoiser within the iterative reconstruction process. Extensive experiments on the FFHQ, CelebA and ImageNet datasets demonstrate that Diff-OneBit gives high-fidelity reconstructed images, outperforming state-of-the-art methods in both reconstruction quality and computational efficiency across 1-bit compressed sensing and logistic regression tasks.
Introduction
In linear measurement models (candes2006robust; donoho2006compressed; foucart2013invitation), the goal is to recover a high-dimensional signal $\mathbf{x}^{*}\in\mathbb{R}^{N}$ from noisy linear measurements:
@@ Introduction
where $\mathbf{A}=\left[\mathbf{a}_{1},\dots,\mathbf{a}_{M}\right]^{\top}\in\mathbb{R}^{M\times N}$ is the forward matrix, $\bm{\epsilon}=[\epsilon_{1},\ldots,\epsilon_{M}]^{\top}\in\mathbb{R}^{M}$ is the noise vector, and $\mathbf{y}=[y_{1},\ldots,y_{M}]^{\top}\in\mathbb{R}^{M}$ represents the acquired measurements. To address this typically ill-posed measurement model, especially in the high-dimensional regime where $M\ll N$, a series of works has developed effective recovery methods by exploiting hand-crafted structural priors of $\mathbf{x}^{*}$, such as sparsity or low-rankness (chen2013low; li2013compressed; nguyen2012robust; nguyen2013exact; xu2012outlier).
However, the assumption of infinite measurement precision in linear measurement models is often unrealistic in hardware implementation. In practice, measurements are quantized to a finite number of bits. In particular, 1-bit quantization is valued for its low hardware cost and robustness to non-linear distortions (boufounos20081; boufounos2010reconstruction). Both 1-bit compressed sensing (CS) and logistic regression are examples of the 1-bit quantized measurement model. For 1-bit CS, measurements are quantized to their signs:
@@
1-bit quantized measurement models have also been widely studied under hand-crafted structural priors such as sparsity and low-rankness (plan2012robust; yan2012robust; plan2013one; jacques2013robust; davenport20141; knudson2016one; plan2017high).
Priors have shifted from hand-crafted designs to leveraging the expressive capabilities of conventional generative models, such as variational autoencoders (VAEs) and generative adversarial networks (GANs) (bora2017compressed; liu2020information). Recently, diffusion models (DMs) (sohl2015deep; ho2020denoising) stand out as powerful generative priors for capturing the statistical properties of natural images. Reconstruction methods using DMs generally fall into two categories: Supervised methods, which train neural networks end-to-end on paired clean and degraded images for specific measurement models, and unsupervised methods, which utilize pretrained unconditional diffusion models and knowledge of the degradation process. Recently, unsupervised methods have been favored for their versatility. They typically integrate a data-fidelity term with a DM-based prior within a gradient based optimization or sampling framework, achieving reliable reconstruction quality across a wide range of linear and non-linear measurement models (kadkhodaie2021stochastic; jalal2021robust; choi2021ilvr; chung2022improving; chung2023diffusion; zhu2023denoising; kawar2022denoising; wang2023zero; song2023pseudoinverse; song2024solving; mardani2024variational; zhang2025improving; zheng2025integrating).
Specifically, under the assumption that the nonlinear forward operator $\mathcal{A}\,:\,\mathbb{R}^{N}\to\mathbb{R}^{M}$ is differentiable, approaches such as DPS (chung2023diffusion), DiffPIR (zhu2023denoising), and DAPS (zhang2025improving) can handle nonlinear measurement models of the form $\mathbf{y}=\mathcal{A}(\mathbf{x}^{*})+\bm{\epsilon}$ through gradient based iterative procedures. However, these methods are not directly applicable to 1-bit CS or logistic regression, as the corresponding forward operators are either non-differentiable or lack an explicit formulation.
Contributions
To address this limitation, we draw inspiration from plug-and-play (PnP) frameworks (venkatakrishnan2013plug; chan2016plug; sreehari2016plug; kamilov2017plug) and propose Diff-OneBit, which is a PnP method designed to tackle the 1-bit CS problem using pretrained DMs. To overcome the challenge posed by the non-differentiable link function in 1-bit CS, our approach decouples the data and prior terms via half-quadratic splitting (HQS) (geman1995nonlinear) and introduces a differentiable surrogate function for the data-fidelity term, solving them iteratively within the diffusion sampling framework. We further adapt Diff-OneBit to address the logistic regression problem. Evaluations of Diff-OneBit on 1-bit CS and logistic regression tasks using the FFHQ and CelebA datasets demonstrate that it achieves superior image reconstruction quality with greater efficiency compared to competing methods.
Related work
@@
Signal recovery with conventional generative models: Building on the pioneering work of (bora2017compressed), prior research has explored signal recovery problems using conventional generative models. Subsequent studies have introduced novel architectures, including untrained deep image priors (van2018compressed), underparameterized deep decoders (heckel2019deep), and invertible networks designed to eliminate representation error (asim2020invertible). Additionally, several methods with provable guarantees have been developed, such as those based on Langevin dynamics (nguyen2021provable) and posterior sampling (jalal2021instance). The application scope of these priors has expanded to address problems like phase retrieval (hand2018phase; liu2021towards; liu2022misspecified), generative principal component analysis (liu2022generative; chen2025solving) and generalized eigenvalue problems (liu2024generalized). This progress has been supported by extensive studies confirming the adversarial robustness of these methods (genzel2022solving) and comprehensive surveys of the field (ongie2020deep; scarlett2023theoretical).
Notably, there have been specific investigations into 1-bit CS using conventional generative models. For instance, qiu2020robust employ ReLU generative networks with dithering techniques to enable signal recovery. liu2020sample establish recovery guarantees for 1-bit CS using conventional generative models. Furthermore, single-index models encompassing 1-bit quantized measurement models have been studied under conventional generative priors (wei2019statistical; liu2020generalized; liu2022projected; liu2022non; chen2023unified). However, these works focus on recovery guarantees related to optimal solutions for empirical risk minimization problems or approaches that approximately optimize over the range of conventional single-step generative models, without considering plug-and-play frameworks.
@@
In QCS-SGM (meng2023quantized), a DM serves as a prior for recovering signals from noisy quantized measurements through posterior sampling, integrating a noise-perturbed pseudo-likelihood score with the score of the DM. However, QCS-SGM relies on annealed Langevin dynamics, which demands a large number of iterations and incurs substantial computational overhead. SIM-DMIS (tang2025learning) investigates single index models that encompass 1-bit quantized measurement models with a diffusion prior. The approach first obtains an appropriate initial estimate, then reconstructs the underlying signal using a single round of unconditional sampling combined with partial inversion of the DM, which is an operation that approximates projecting the initial estimate onto the range the DM.
Background
Diffusion models
DMs are generative models that map data to noise through a pre-defined forward diffusion process (ho2020denoising; karras2022elucidating; song2019generative; song2020improved; song2021score). Specifically, the forward process perturbs a data sample $\mathbf{x}_{0}\sim p_{\text{data}}(\mathbf{x})$ over a continuous time variable $t\in[0,T]$ according to a Gaussian transition kernel:
@@
To generate a sample, DMs learn to reverse this process, starting from pure Gaussian noise $\mathbf{x}_{T}\sim\mathcal{N}(\bm{0},\mathbf{I})$ and iteratively denoising it to produce a sample from the data distribution. Specifically, a neural network $\bm{\epsilon}_{\mathbf{\theta}}(\mathbf{x}_{t},t)$ is trained to predict the noise component $\bm{\epsilon}$ from the noisy state $\mathbf{x}_{t}=\alpha_{t}\mathbf{x}_{0}+\sigma_{t}\bm{\epsilon}$ by minimizing the following objective (song2019generative; song2021denoising):
$\mathbb{E}_{\mathbf{x}_{0}\sim p_{0},\bm{\epsilon}\sim\mathcal{N}(\bm{0},\bm{I}),t\sim\mathcal{U}(0,T)}[\|\bm{\epsilon}_{\mathbf{\theta}}(\mathbf{x}_{t},t)-\bm{\epsilon}\|_{2}^{2}].$ (5)
The pretrained noise predictor $\bm{\epsilon}_{\bm{\theta}}(\mathbf{x}_{t},t)$ provides an estimate of the score function via the relation $\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t})\approx-\bm{\epsilon}_{\bm{\theta}}(\mathbf{x}_{t},t)/\sigma_{t}$. Thus, sampling from DMs can be performed by numerically solving a stochastic differential equation (SDE) (song2021score):
$\mathrm{d}\mathbf{x}_{t}=\left[f(t)\mathbf{x}_{t}+\frac{g^{2}(t)}{\sigma_{t}}\bm{\epsilon}_{\bm{\theta}}(\mathbf{x}_{t},t)\right]\mathrm{d}t+g(t)\mathrm{d}\mathbf{w}_{t},$ (6)
@@
Conditional diffusion models for signal recovery
@@
For a linear measurement model with additive Gaussian noise $\bm{\epsilon}\sim\mathcal{N}(\mathbf{0},\sigma^{2}\mathbf{I})$ as in Eq. (1), the conditional score is directly related to the conditional expectation as follows (daras2024survey):
@@
Here, $\mu_{t}:=\lambda(\alpha_{t}^{2}/\sigma_{t}^{2})$ with $\lambda>0$ being a hyperparameter is a penalty parameter and $\mathbb{E}[\mathbf{x}_{0}|\mathbf{x}_{t}]$ can be directly obtained using Tweedie’s formula (daras2024survey). Alternatively, approaches such as DPS approximate $\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{y}|\mathbf{x}_{t})$ using Bayes’ rule with $\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t}|\mathbf{y})=\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{y}|\mathbf{x}_{t})+\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t})$, where the term $\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t})$ is simply the unconditional score function. However, these frameworks cannot be directly extended to handle 1-bit quantized measurement models for which both $\mathbb{E}[\mathbf{x}_{0}|\mathbf{x}_{t},\mathbf{y}]$ and $\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{y}|\mathbf{x}_{t})$ are difficult to approximate.
Method
@@
$=\arg\min_{\mathbf{x}}\underbrace{-\log p(\mathbf{y}|\mathbf{x})}_{\text{Data Fidelity Term}}+\underbrace{(-\log p(\mathbf{x}))}_{\text{Prior Regularization Term}}.$ (12)
@@
Combining these two cases into a single one, yielding:
@@
where $\mu>0$ is the penalty parameter and $\hat{\mathbf{z}}$ is an estimated vector. The differentiability of $\mathcal{L}(\cdot;\mathbf{y})$ enables efficient optimization using gradient based methods such as stochastic gradient descent or Adam optimizer (kingma2014adam). The $\mathbf{z}$-update rule involves solving an optimization problem of the following form:
@@
The optimization problem in Eq. (26) can be thought of as a denoising problem, and its optimal solution can be approximated using a DM through the single-step Tweedie’s formula, see, e.g.,  (zhu2023denoising; venkatakrishnan2013plug), for which we provide a detailed analysis in the appendix.
@@
Adaption for logistic regression
@@
$=-\sum_{i=1}^{M}\left[\frac{1+y_{i}}{2}\log\left(\frac{1}{2}+\frac{1}{2}\tanh\left(\frac{\mathbf{a}_{i}^{\top}\mathbf{x}}{2}\right)\right)\right$.
$\quad\left.+\frac{1-y_{i}}{2}\log\left(\frac{1}{2}-\frac{1}{2}\tanh\left(\frac{\mathbf{a}_{i}^{\top}\mathbf{x}}{2}\right)\right)\right].$ (31)
@@
1:Measurements $\mathbf{y}$, forward matrix $\mathbf{A}$, pretrained noise prediction network $\bm{\epsilon_{\theta}}$, maximum number of iterations $K$, data-fidelity function $\mathcal{L}(\cdot;\mathbf{y})$, schedules $\{\alpha_{t_{k}},\sigma_{t_{k}}\}_{k=0}^{K}$, penalty coefficient $\lambda$, pre-calculated $\mu_{k}=\lambda(\alpha_{t_{k}}^{2}/\sigma_{t_{k}}^{2})$
@@
5:  $\mathbf{\tilde{z}}_{t_{k}}\leftarrow(\mathbf{\tilde{x}}_{t_{k}}-\sigma_{t_{k}}\bm{\epsilon}_{\theta}(\mathbf{\tilde{x}}_{t_{k}},t_{k}))/\alpha_{t_{k}}$

### v2

@@ Abstract
Diffusion models (DMs) have demonstrated to be powerful priors for signal recovery, but their application to 1-bit quantization tasks, such as 1-bit compressed sensing and logistic regression, remains a challenge. This difficulty stems from the inherent non-linear link function in these tasks, which is either non-differentiable or lacks an explicit characterization. To tackle this issue, we introduce Diff-OneBit, which is a fast and effective DM-based approach for signal recovery under 1-bit quantization. Diff-OneBit addresses the challenge posed by non-differentiable or implicit links functions via leveraging a differentiable surrogate likelihood function to model 1-bit quantization, thereby enabling gradient based iterations. This function is integrated into a flexible plug-and-play framework that decouples the data-fidelity term from the diffusion prior, allowing any pretrained DM to act as a denoiser within the iterative reconstruction process. Extensive experiments on the FFHQ, CelebA and ImageNet datasets demonstrate that Diff-OneBit gives high-fidelity reconstructed images, outperforming state-of-the-art methods in both reconstruction quality and computational efficiency across 1-bit compressed sensing and logistic regression tasks. Our code is available at https://github.com/Chenyouming123/DiffOneBit.
I Introduction
In linear measurement models [5, 17, 18], the goal is to recover a high-dimensional signal $\mathbf{x}^{*}\in\mathbb{R}^{N}$ from noisy linear measurements:
@@ I Introduction
where $\mathbf{A}=\left[\mathbf{a}_{1},\dots,\mathbf{a}_{M}\right]^{\top}\in\mathbb{R}^{M\times N}$ is the forward matrix, $\bm{\epsilon}=[\epsilon_{1},\ldots,\epsilon_{M}]^{\top}\in\mathbb{R}^{M}$ is the noise vector, and $\mathbf{y}=[y_{1},\ldots,y_{M}]^{\top}\in\mathbb{R}^{M}$ represents the acquired measurements. To address this typically ill-posed measurement model, especially in the high-dimensional regime where $M\ll N$, a series of works has developed effective recovery methods by exploiting hand-crafted structural priors of $\mathbf{x}^{*}$, such as sparsity or low-rankness [9, 36, 48, 49, 70].
However, the assumption of infinite measurement precision in linear measurement models is often unrealistic in hardware implementation. In practice, measurements are quantized to a finite number of bits. In particular, 1-bit quantization is valued for its low hardware cost and robustness to non-linear distortions [3, 4]. Both 1-bit compressed sensing (CS) and logistic regression are examples of the 1-bit quantized measurement model. For 1-bit CS, measurements are quantized to their signs:
@@
1-bit quantized measurement models have also been widely studied under hand-crafted structural priors such as sparsity and low-rankness [53, 71, 54, 25, 14, 35, 52].
Priors have shifted from hand-crafted designs to leveraging the expressive capabilities of conventional generative models, such as variational autoencoders (VAEs) and generative adversarial networks (GANs) [2, 43]. Recently, diffusion models (DMs) [57, 24] stand out as powerful generative priors for capturing the statistical properties of natural images. Reconstruction methods using DMs generally fall into two categories: Supervised methods, which train neural networks end-to-end on paired clean and degraded images for specific measurement models, and unsupervised methods, which utilize pretrained unconditional diffusion models and knowledge of the degradation process. Recently, unsupervised methods have been favored for their versatility. They typically integrate a data-fidelity term with a DM-based prior within a gradient based optimization or sampling framework, achieving reliable reconstruction quality across a wide range of linear and non-linear measurement models [28, 27, 10, 12, 11, 75, 33, 68, 60, 58, 46, 72, 74].
Specifically, under the assumption that the nonlinear forward operator $\mathcal{A}\,:\,\mathbb{R}^{N}\to\mathbb{R}^{M}$ is differentiable, approaches such as DPS [11], DiffPIR [75], and DAPS [72] can handle nonlinear measurement models of the form $\mathbf{y}=\mathcal{A}(\mathbf{x}^{*})+\bm{\epsilon}$ through gradient based iterative procedures. However, these methods are not directly applicable to 1-bit CS or logistic regression, as the corresponding forward operators are either non-differentiable or lack an explicit formulation.
Contributions: To address this limitation, we draw inspiration from plug-and-play (PnP) frameworks [67, 6, 64, 29] and propose Diff-OneBit, which is a PnP method designed to tackle the 1-bit CS problem using pretrained DMs. To overcome the challenge posed by the non-differentiable link function in 1-bit CS, our approach decouples the data and prior terms via half-quadratic splitting (HQS) [19] and introduces a differentiable surrogate function for the data-fidelity term, solving them iteratively within the diffusion sampling framework. We further adapt Diff-OneBit to address the logistic regression problem. Evaluations of Diff-OneBit on 1-bit CS and logistic regression tasks using the FFHQ and CelebA datasets demonstrate that it achieves superior image reconstruction quality with greater efficiency compared to competing methods.
I-A Related work
@@
Signal recovery with conventional generative models: Building on the pioneering work of [2], prior research has explored signal recovery problems using conventional generative models. Subsequent studies have introduced novel architectures, including untrained deep image priors [66], underparameterized deep decoders [22], and invertible networks designed to eliminate representation error [1]. Additionally, several methods with provable guarantees have been developed, such as those based on Langevin dynamics [50] and posterior sampling [26]. The application scope of these priors has expanded to address problems like phase retrieval [21, 38, 45], generative principal component analysis [42, 7] and generalized eigenvalue problems [41]. This progress has been supported by extensive studies confirming the adversarial robustness of these methods [20] and comprehensive surveys of the field [51, 56].
Notably, there have been specific investigations into 1-bit CS using conventional generative models. For instance, the work [55] employ ReLU generative networks with dithering techniques to enable signal recovery. The work [39] establish recovery guarantees for 1-bit CS using conventional generative models. Furthermore, single-index models encompassing 1-bit quantized measurement models have been studied under conventional generative priors [69, 44, 40, 37, 8]. However, these works focus on recovery guarantees related to optimal solutions for empirical risk minimization problems or approaches that approximately optimize over the range of conventional single-step generative models, without considering plug-and-play frameworks.
@@
In QCS-SGM [47], a DM serves as a prior for recovering signals from noisy quantized measurements through posterior sampling, integrating a noise-perturbed pseudo-likelihood score with the score of the DM. However, QCS-SGM relies on annealed Langevin dynamics, which demands a large number of iterations and incurs substantial computational overhead. SIM-DMIS [65] investigates single index models that encompass 1-bit quantized measurement models with a diffusion prior. The approach first obtains an appropriate initial estimate, then reconstructs the underlying signal using a single round of unconditional sampling combined with partial inversion of the DM, which is an operation that approximates projecting the initial estimate onto the range the DM.
II Background
II-A Diffusion models
DMs are generative models that map data to noise through a pre-defined forward diffusion process [24, 31, 61, 62, 63]. Specifically, the forward process perturbs a data sample $\mathbf{x}_{0}\sim p_{\text{data}}$ over a continuous time variable $t\in[0,T]$ according to a Gaussian transition kernel:
@@ II-A Diffusion models
To generate a sample, DMs learn to reverse this process, starting from pure Gaussian noise $\mathbf{x}_{T}\sim\mathcal{N}(\bm{0},\mathbf{I})$ and iteratively denoising it to produce a sample from the data distribution. Specifically, a neural network $\bm{\epsilon}_{\mathbf{\bm{\theta}}}(\mathbf{x}_{t},t)$ is trained to predict the noise component $\bm{\epsilon}$ from the noisy state $\mathbf{x}_{t}=\alpha_{t}\mathbf{x}_{0}+\sigma_{t}\bm{\epsilon}$ by minimizing the following objective [61, 59]:
$\mathbb{E}_{\mathbf{x}_{0}\sim p_{0},\bm{\epsilon}\sim\mathcal{N}(\bm{0},\bm{I}),t\sim\mathcal{U}(0,T)}[\|\bm{\epsilon}_{\mathbf{\bm{\theta}}}(\mathbf{x}_{t},t)-\bm{\epsilon}\|_{2}^{2}].$ (5)
The pretrained noise predictor $\bm{\epsilon}_{\bm{\theta}}(\mathbf{x}_{t},t)$ provides an estimate of the score function via the relation $\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t})\approx-\bm{\epsilon}_{\bm{\theta}}(\mathbf{x}_{t},t)/\sigma_{t}$. Thus, sampling from DMs can be performed by numerically solving a stochastic differential equation (SDE) [63]:
$\mathrm{d}\mathbf{x}_{t}=\left(f(t)\mathbf{x}_{t}+\frac{g^{2}(t)}{\sigma_{t}}\bm{\epsilon}_{\bm{\theta}}(\mathbf{x}_{t},t)\right)\mathrm{d}t+g(t)\mathrm{d}\mathbf{w}_{t},$ (6)
@@
II-B Conditional diffusion models for signal recovery
@@
For a linear measurement model with additive Gaussian noise $\bm{\epsilon}\sim\mathcal{N}(\mathbf{0},\sigma^{2}\mathbf{I})$ as in Eq. (1), the conditional score is directly related to the conditional expectation as follows [13]:
@@
Here, $\mu_{t}:=\lambda(\alpha_{t}^{2}/\sigma_{t}^{2})$ with $\lambda>0$ being a hyperparameter and $\mathbb{E}[\mathbf{x}_{0}|\mathbf{x}_{t}]$ can be directly obtained using Tweedie’s formula [13]. Alternatively, approaches such as DPS approximate $\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{y}|\mathbf{x}_{t})$ using Bayes’ rule with
$\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t}|\mathbf{y})=\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{y}|\mathbf{x}_{t})+\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t}),$ (11)
where the term $\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t})$ is simply the unconditional score function. However, these frameworks cannot be directly extended to handle 1-bit quantized measurement models for which both $\mathbb{E}[\mathbf{x}_{0}|\mathbf{x}_{t},\mathbf{y}]$ and $\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{y}|\mathbf{x}_{t})$ are difficult to approximate.
III Method
@@
$=\arg\min_{\mathbf{x}}\underbrace{-\log p(\mathbf{y}|\mathbf{x})}_{\text{Data Fidelity Term}}\quad+\underbrace{(-\log p(\mathbf{x})).}_{\text{Prior Regularization Term}}$ (13)
@@ III-A Approach for 1-bit CS
Combining these two cases into a single one, yielding
@@
where $\mu>0$ is the penalty parameter and $\hat{\mathbf{z}}$ is an estimated vector. The differentiability of $\mathcal{L}(\cdot;\mathbf{y})$ enables efficient optimization using gradient based methods such as stochastic gradient descent or Adam optimizer [34]. The $\mathbf{z}$-update rule involves solving an optimization problem of the following form:
@@
The optimization problem in Eq. (27) can be thought of as a denoising problem, and its optimal solution can be approximated using a DM through the single-step Tweedie’s formula, see, e.g., [75, 67], for which we provide a detailed analysis in the appendix.
@@
III-B Adaption for logistic regression
@@
$=-\sum_{i=1}^{M}\left(\frac{1+y_{i}}{2}\log\left(\frac{1}{2}+\frac{1}{2}\tanh\left(\frac{\mathbf{a}_{i}^{\top}\mathbf{x}}{2}\right)\right)+\frac{1-y_{i}}{2}\log\left(\frac{1}{2}-\frac{1}{2}\tanh\left(\frac{\mathbf{a}_{i}^{\top}\mathbf{x}}{2}\right)\right)\right).$ (32)
@@
1:Measurements $\mathbf{y}$, forward matrix $\mathbf{A}$, pretrained noise prediction network $\bm{\epsilon_{\bm{\theta}}}$, maximum number of iterations $K$, data-fidelity function $\mathcal{L}(\cdot;\mathbf{y})$, schedules $\{\alpha_{t_{k}},\sigma_{t_{k}}\}_{k=0}^{K}$, penalty coefficient $\lambda$, pre-calculated $\mu_{k}=\lambda(\alpha_{t_{k}}^{2}/\sigma_{t_{k}}^{2})$
@@
5:  $\mathbf{\tilde{z}}_{t_{k}}\leftarrow(\mathbf{\tilde{x}}_{t_{k}}-\sigma_{t_{k}}\bm{\epsilon}_{\bm{\theta}}(\mathbf{\tilde{x}}_{t_{k}},t_{k}))/\alpha_{t_{k}}$